OpenStudy (anonymous):

1 MEDAL HELP!!!! When planning his crops, Farmer Tim knows that his 7 acres can support peas and carrots. He wants to make $7,000 from his crops. He can make $2,000 per acre of peas (the variable p) and $1,000 per acre of carrots (the variable c). Which equation below would be a constraint in his system of equations?

OpenStudy (anonymous):

p + c = 7

OpenStudy (anonymous):

aprilann143 can u help me? i tagged you

OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 36 y = x - 6

OpenStudy (anonymous):

how do you do this one?

OpenStudy (anonymous):

by any chance did u get the answer to this one ?